OpenStudy (lilblister):

Is 2x^2+8x+4!a perfect square. Yes or no? I think it's no?

OpenStudy (ness9630):

Have you tried factoring a constant out first?

OpenStudy (lilblister):

Well arent you supposed to try squaring the first and last numbers?

OpenStudy (ness9630):

Yes, but if you factor it might be easier: \(\Large 2x^2+8x+4 = 2(x^2+4x+2 )\) Does the inside look like a perfect square?

OpenStudy (lilblister):

Wait no ? Idk lol im confused

OpenStudy (ness9630):

And actually yes, you were right, squaring the last and first terms would help. Do know what number, if square, gives you 2 and 4?

OpenStudy (lilblister):

16 squared gives you 4.

OpenStudy (lilblister):

And 4 squared givse you two.

OpenStudy (ness9630):

The square root of those numbers, yes. Squared tho, only \(\Large \sqrt{2} \) and 2 give you 4 and 2 when squared.

OpenStudy (lilblister):

Oh lol sorry i get those two mixed up sometimes XD

OpenStudy (lilblister):

So it is a perfect square then?

OpenStudy (lilblister):

be right back

OpenStudy (lilblister):

ok im back c:

OpenStudy (ness9630):

Let's check, a perfect square looks something like \( (a+b)^2 \) or \( (a-b)^2 \) Since there are no minuses, we'll use the first one and plug in those numbers: \(\Large [\sqrt{2}x+2 ]^2 \) And alrightt

OpenStudy (lilblister):

So it is then?

OpenStudy (ness9630):

If you finish it, you'll get \(\Large 2x^2+4\sqrt{2}x+4 \).. does tha look like the previous problem? :P

OpenStudy (lilblister):

No?

OpenStudy (lilblister):

lol. Im sorry. Math confuses me XD

OpenStudy (lilblister):

@Ness9630

OpenStudy (ness9630):

It's not, and sorry, I probably should have done this earlier, but if it was a perfect square, it would only give one root, whereas this one gives us two. So you're right, no :P Do you understand what I mean?

OpenStudy (lilblister):

Yes i do! Thank you so much c: Your very helpful! Thank you for your time

OpenStudy (ness9630):

Alright, and you're welcome \(\LARGE \color{red}{\star^{\color{green}{\star}}} \)